### v1

@@ 6 Shower generation
To simulate extensive air showers triggered by UHE particles and the radio emission from them GRANDlib relies on either ZHAireS [15] or CoREAS [16], Monte-Carlo shower simulators that are run separately to GRANDlib, and whose output is fed to it. For details, we defer to Refs. [15, 16]; below, we merely sketch the process of generating shower simulations. ZHAireS and CoREAS yield similar results [32, 33, 34] and have been shown to match radio measurements [35, 36, 37].
To generate a shower simulation, the user specifies a few input parameters that define the shower, i.e., the primary particle type, its initial energy, arrival direction—defined by the zenith angle, $\theta$, and the azimuth angle, $\phi$—and a hadronic interaction model with which the particles will interact in the atmosphere (QGSJET-II-04 [38] for the results in this paper). The environment in which the shower propagates is characterized by specifying the geomagnetic field amplitude, $B$, its inclination $i=\arctan{(B_{z}/B_{x})}$ at the shower location, where $B_{x}$ and $B_{z}$ are the field components along the x- and z-axes, and a model of the matter density profile of the atmosphere, e.g., the US atmospheric model [39] or a model generated using the Global Data Assimilation System (GDAS). The simulations used in this paper were generated using an atmospheric model based on GDAS data for a location near the GRANDProto300 site. The user also provides a list of the antenna positions on the ground at which the electric field of the radio emission from the shower should be computed.
@@ 7 Radio antenna response
In GRANDlib, the effective length of each of the antenna arms of the GRAND HorizonAntenna [1]—designed to be sensitive to very inclined air showers—is stored as a pre-computed table for the X-, Y-, and Z-arms, for $f\in[30,250]$ MHz, $\theta\in[0^{\circ},90^{\circ}]$, and $\phi\in[0^{\circ},360^{\circ}]$, and is interpolated for values that are not pre-computed, its real and imaginary parts treated independently. The effective length accounts for reflections of the radio signals off of the ground, and for the height of the antenna pole, of 3.2 m in the HorizonAntenna.
Figure 5 shows the magnitude of the effective length for the X- and Z-arms of a HorizonAntenna at a frequency of 150 MHz, for an incoming electric field along different directions. The effective length is maximum when a component of the electric field oscillates parallel to the antenna arm. In Fig. 5, for the X-arm (and also for the Y-arm, not shown), the $\phi$-polarization of the electric field oscillates parallel to the X-arm of the antenna when the shower direction is $\phi=90^{\circ}$ or $\phi=270^{\circ}$. This results in the effective length being maximum along the $\phi$-polarization where it is minimum for the $\theta$-polarization. For the Z-arm, because it is vertical (Fig. 4), its effective length is almost exclusively in the $\theta$-polarization.
@@ 8 Signal processing
First, we use ZHAireS or CoREAS to simulate the electric fields of radio signals generated in extensive air showers. Second, at each DU, the electric field is detected by a passive antenna and converted into an open-circuit voltage signal using the antenna effective length (Section 7). We then add noise—of Galactic and anthropogenic origin—to the open-circuit voltage. Finally, the open-circuit voltage is transmitted through various components of the RF chain (Section 8.3) of the DU to generate a final output voltage, which is then digitized by an analog-to-digital converter (ADC). By default, digitization is carried out with 14-bit precision at a sampling rate of 500 million samples per second and the conversion factor is 1 ADC equivalent to 109.86 $\mu$[11]; these values can be adjusted in the code.
@@ 8.2 Galactic noise
To compute the Galactic radio noise, GRANDlib uses LFMap [48], which returns the brightness temperature across the sky. The response of a GRAND HorizonAntenna to Galactic noise is pre-computed and tabulated for frequencies of 30–250 MHz and local sidereal times (LSTs) of 0–24 hours. Using it, the galaxy module returns the voltage induced by the Galactic noise induced at each antenna arm, at requested values of frequency and LST.
@@ 8.5 Run-time benchmarks
On average, the time required by GRANDlib to convert the electric field of a shower into an output voltage in all of the 300 DUs that make up GRANDProto300 is about 9 s, when running it on a single core of a personal computer with a 1.3-GHz Intel Core i5 processor.
@@ 11 Summary and outlook
GRANDlib is in active development by the GRAND Collaboration to meet the evolving needs of the experiment. Planned improvements include a smoother interface with ZHAireS and CoREAS, more advanced visualization tools, and tools to streamline the processing pipeline and data analysis. There is also ongoing work in event reconstruction, i.e., inferring properties such as the shower energy and direction, and the identity of the primary particle, from the detected radio signals.
@@ 12 Acknowledgements
The GRAND Collaboration is grateful to Arno Engels, from Radboud University Nijmegen, for providing the diagram of the GRAND HorizonAntenna in Fig. 4, and to Fei Gao, from Xidian University, for their work on antenna design. The GRAND Collaboration acknowledges the support from the following funding agencies and grants. Brazil: Conselho Nacional de Desenvolvimento Cientiífico e Tecnológico (CNPq); Fundação de Amparo à Pesquisa do Estado de Rio de Janeiro (FAPERJ); Coordenação Aperfeiçoamento de Pessoal de Nível Superior (CAPES). China: NAOC, National SKA Program of China (Grant No. 2020SKA0110200); Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061); Program for Innovative Talents and Entrepreneur in Jiangsu, and High-end Foreign Expert Introduction Program in China (No. G2023061006L); China Scholarship Council (No. 202306010363). Denmark: Villum Fonden (project no. 29388). France: “Emergences” Program of Sorbonne Université; France-China Particle Physics Laboratory; Program National des Hautes Energies of INSU; for IAP—Agence Nationale de la Recherche (“APACHE” ANR-16-CE31-0001, “NUTRIG” ANR-21-CE31-0025, ANR-23-CPJ1-0103-01), CNRS Programme IEA Argentine (“ASTRONU”, 303475), CNRS Programme Blanc MITI (“GRAND” 2023.1 268448), CNRS Programme AMORCE (“GRAND” 258540); Fulbright-France Programme; IAP+LPNHE—Programme National des Hautes Energies of CNRS/INSU with INP and IN2P3, co-funded by CEA and CNES; IAP+LPNHE+KIT—NuTRIG project, Agence Nationale de la Recherche (ANR-21-CE31-0025); IAP+VUB: PHC TOURNESOL program 48705Z. Germany: NuTRIG project, Deutsche Forschungsgemeinschaft (DFG, Projektnummer 490843803); Helmholtz—OCPC Postdoc-Program. Poland: Polish National Agency for Academic Exchange within Polish Returns Program no. PPN/PPO/2020/1/00024/U/00001,174; NCN OPUS grant no. 2022/45/B/ST2/0288. USA: U.S. National Science Foundation under Grant No. 2418730. Computer simulations were performed using computing resources at the CCIN2P3 Computing Centre (Lyon/Villeurbanne, France), partnership between CNRS/IN2P3 and CEA/DSM/Irfu, and computing resources supported by the Chinese Academy of Sciences.
@@ Appendix D GRANDlib package dependencies
black jupyter-notebook numpy SonarQube
cffi lxml paramiko sshtunnel
coverage matplotlib psycopg2 typing_extensions
docker mypy pylint GULL, TURTLE

### v2

@@ 6 Shower generation
To simulate extensive air showers triggered by UHE particles and the radio emission from them, GRANDlib relies on either ZHAireS [17] or CoREAS [18], Monte-Carlo shower simulators that are run separately to GRANDlib, and whose output is fed to it. For details, we defer to Refs. [17, 18]; below, we merely sketch the process of generating shower simulations. ZHAireS and CoREAS yield similar results [37, 38, 39] and have been shown to match radio measurements [40, 41, 42].
To generate a shower simulation, the user specifies a few input parameters that define the shower, i.e., the primary particle type, its initial energy, arrival direction—defined by the zenith angle, $\theta$, and the azimuth angle, $\phi$—and a hadronic interaction model with which the particles will interact in the atmosphere (QGSJET-II-04 [43] for the results in this paper). The environment in which the shower propagates is characterized by specifying the geomagnetic field amplitude, $B$, its inclination, $i=\arctan{(B_{z}/B_{x})}$, at the shower location, where $B_{x}$ and $B_{z}$ are the field components along the x- and z-axes, and a model of the matter density profile of the atmosphere, e.g., the US atmospheric model [44] or a model generated using the Global Data Assimilation System (GDAS). The simulations used in this paper were generated using an atmospheric model based on GDAS data for a location near the GRANDProto300 site. The user also provides a list of the antenna positions on the ground at which the electric field of the radio emission from the shower should be computed.
@@ 7 Radio antenna response
In GRANDlib, the effective length of each of the antenna arms of the GRAND HorizonAntenna [1]—designed to be sensitive to very inclined air showers—is stored as a pre-computed table for the X-, Y-, and Z-arms, for $f\in[30,250]$ MHz, $\theta\in[0^{\circ},90^{\circ}]$, and $\phi\in[0^{\circ},360^{\circ}]$, and is interpolated for values that are not pre-computed, its real and imaginary parts treated independently. The effective length accounts for reflections of the radio signals off of the ground, and for the height of the antenna pole, of 3.2 m in the HorizonAntenna. The effective length used in GRANDlib holds regardless of whether the radio signal is short and broadband (i.e., wide in its frequency range), or long and narrowband.
Figure 5 shows the magnitude of the effective length for the X- and Z-arms of a HorizonAntenna at a frequency of 150 MHz, for an incoming electric field along different directions. This frequency is representative of the shape of the structures in the effective length, but the structures change in size and number for other frequencies; this is accounted for in GRANDlib. The effective length is maximum when a component of the electric field oscillates parallel to the antenna arm.
In Fig. 5, for the X-arm (and also for the Y-arm, not shown), the $\phi$-polarization of the electric field oscillates parallel to the X-arm of the antenna when the shower direction is $\phi=90^{\circ}$ or $\phi=270^{\circ}$. This results in the effective length being maximum along the $\phi$-polarization where it is minimum for the $\theta$-polarization. For the Z-arm, because it is vertical (Fig. 4), its effective length is almost exclusively in the $\theta$-polarization. The striped pattern in the $\theta$-polarization is due to the destructive interference of the radio signal with itself after reflecting off the ground. The reflection lengthens the path of the signal, delaying its arrival to the antenna by up to tens of ns, with the exact delay being different for different frequencies (and for different incidence angles of the signal), and the net delay being due to the interference of the different frequencies. A detailed study of the antenna effective length will be presented elsewhere.
@@ 8 Signal processing
First, we use ZHAireS or CoREAS to simulate the electric fields of radio signals generated in extensive air showers. Second, at each DU, the electric field is detected by a passive antenna and converted into an open-circuit voltage signal using the antenna effective length (Section 7). We then add noise—of Galactic and anthropogenic origin—to the open-circuit voltage. Finally, the open-circuit voltage is transmitted through various components of the RF chain (Section 8.3) of the DU to generate a final output voltage, which is then digitized by an analog-to-digital converter (ADC). By default, digitization is carried out with 14-bit precision at a sampling rate of 500 million samples per second and the conversion factor is 1 ADC equivalent to 109.86 $\mu$[12]; these values can be adjusted in the code. The digitization stores traces from four channels, each with a trace length of 4096 bins of 2 ns, with each bin requiring 2 bytes of storage. Three channels store data from the X-, Y-, and Z-arms of the antenna; the fourth channel typically collects noise, and is used for testing and as backup.
@@ 8.2 Galactic noise
To compute the Galactic radio noise, GRANDlib uses LFMap [52], which returns the brightness temperature across the sky, which included the cosmic microwave background, an isotropic component, and an anisotropic component from the Milky Way. The response of a GRAND HorizonAntenna to Galactic noise is pre-computed and tabulated for frequencies of 30–250 MHz and local sidereal times (LSTs) of 0–24 hours. Using it, the galaxy module returns the voltage induced by the Galactic noise induced at each antenna arm, at requested values of frequency and LST.
GRANDlib does not account for the spatial coherence of the Galactic noise registered by neighboring DUs throughout the array, which is expected to be low due to the sparsity of the array. Instead, the galaxy module mimics the variability of the Galactic noise by multiplying, for each DU, the sky-averaged value of the noise by a different random value of its phase for each antenna arm and at each frequency.
@@ 8.5 Run-time benchmarks
On average, the time required by GRANDlib to convert the electric field of a shower into an output voltage in all the DUs that make up GRANDProto300 is about 13 s, when running it on a single core of a personal computer with a 1.3-GHz Intel Core i5 processor. This run time is small compared to the time required to simulate an extensive air shower and the radio emission from it, which can take several hours.
@@ 11 Summary and outlook
GRANDlib is in active development by the GRAND Collaboration to meet the evolving needs of the experiment. Planned improvements include a smoother interface with ZHAireS and CoREAS, more advanced visualization tools, and tools to streamline the processing pipeline and data analysis. There is also ongoing work in event reconstruction [57, 58], i.e., inferring properties such as the shower energy and direction, and the identity of the primary particle, from the detected radio signals.
@@ 12 Acknowledgements
The GRAND Collaboration is grateful to Arno Engels, from Radboud University Nijmegen, for providing the diagram of the GRAND HorizonAntenna in Fig. 4, and to Fei Gao, from Xidian University, for their work on antenna design. The GRAND Collaboration acknowledges the support from the following funding agencies and grants. Brazil: Conselho Nacional de Desenvolvimento Cientiífico e Tecnológico (CNPq); Fundação de Amparo à Pesquisa do Estado de Rio de Janeiro (FAPERJ); Coordenação Aperfeiçoamento de Pessoal de Nível Superior (CAPES). China: NAOC, National SKA Program of China (Grant No. 2020SKA0110200); Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061); Program for Innovative Talents and Entrepreneurs in Jiangsu, and High-end Foreign Expert Introduction Program in China (No. G2023061006L); China Scholarship Council (No. 202306010363). Denmark: Villum Fonden (project no. 29388). France: “Emergences” Programme of Sorbonne Université; France-China Particle Physics Laboratory; Programme National des Hautes Energies of INSU; for IAP—Agence Nationale de la Recherche (“APACHE” ANR-16-CE31-0001, “NUTRIG” ANR-21-CE31-0025, ANR-23-CPJ1-0103-01), CNRS Programme IEA Argentine (“ASTRONU”, 303475), CNRS Programme Blanc MITI (“GRAND” 2023.1 268448), CNRS Programme AMORCE (“GRAND” 258540); Fulbright-France Programme; IAP+LPNHE—Programme National des Hautes Energies of CNRS/INSU with INP and IN2P3, co-funded by CEA and CNES; IAP+LPNHE+KIT—NuTRIG project, Agence Nationale de la Recherche (ANR-21-CE31-0025); IAP+VUB: PHC TOURNESOL programme 48705Z. Germany: NuTRIG project, Deutsche Forschungsgemeinschaft (DFG, Projektnummer 490843803); Helmholtz—OCPC Postdoc-Program. Poland: Polish National Agency for Academic Exchange within Polish Returns Program no. PPN/PPO/2020/1/00024/U/00001,174; National Science Centre Poland for NCN OPUS grant no. 2022/45/B/ST2/0288. USA: U.S. National Science Foundation under Grant No. 2418730. Computer simulations were performed using computing resources at the CCIN2P3 Computing Centre (Lyon/Villeurbanne, France), partnership between CNRS/IN2P3 and CEA/DSM/Irfu, and computing resources supported by the Chinese Academy of Sciences.
@@ Appendix D GRANDlib package dependencies
black jupyter-notebook paramiko SonarQube
cffi lxml psycopg2 sshtunnel
coverage matplotlib pylint TURTLE
docker mypy ROOT typing_extensions
GULL numpy